OpenStudy (anonymous):

quadratic formula for y = x^2 + 10x + 16

OpenStudy (superhelp101):

do you know the quadratic formula ? ;)

OpenStudy (anonymous):

no need to graph the result

OpenStudy (superhelp101):

are the answer choices as graphs ?

OpenStudy (anonymous):

yes

OpenStudy (superhelp101):

okay well then yes ! but you have to use the quadratic formula for this

OpenStudy (superhelp101):

that is the formula now plug in the numbers from you equation to get the answer

OpenStudy (superhelp101):

1. x^2 -2x - 8x +16 = 0 2. x ( x-2) - 8 ( x- 2) = 0 3. ( x-8) ( x-2) = 0 4. x = 8 OR 2 or do it like this

OpenStudy (anonymous):

this is what it should look like

OpenStudy (anonymous):

Thanks!!!

OpenStudy (superhelp101):

and if you want to graph it then you have to plug the equation into you calculator

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

y = x^2 - 10x + 25 Please help

OpenStudy (anonymous):

your graph should look like this

OpenStudy (anonymous):

Thanks Katie!!!

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

solve y=x^2 - 11x + 30 when x is 3 4 5 6 Y using the quadratic fomula

OpenStudy (anonymous):

when x is 3,4,5 and 6?

OpenStudy (anonymous):

yes